# A Case Study on Energy-Efficient Edge AI Crack Segmentation

### 15 Apr 2026 v1

@@ Abstract
Crack segmentation on edge devices can support continuous infrastructure monitoring and maintenance and thereby help to preserve public safety. Furthermore, autonomous infrastructure monitoring by using Unmanned Aerial Vehicles (UAVs) can reduce inspection risks, as human operators no longer need to enter hazardous areas. Edge processing reduces the cost of inspection by eliminating the need for high resolution image storage for offline processing and mitigates the security risks and bandwidth requirements of streaming to cloud servers. Edge inference is difficult due to the limited memory and computational capabilities of edge devices, which can affect both accuracy and latency. Furthermore, battery-powered devices are subject to strict power and energy constraints. Together, these limitations impose restrictions on the model size and computational complexity that can be deployed close to the sensor. In recent years, Transformers have achieved state-of-the-art accuracy in a variety of applications, including semantic segmentation. However, Transformer-based models are typically large and computationally intensive, making efficient edge deployment difficult. To address this, we first apply knowledge distillation to enhance the performance of the base models. We then use Post-Training Quantization (PTQ) to compress the models further. Additionally, we consider the deployment of these models across multiple edge platforms. To maximize energy efficiency, we design and implement a custom hardware architecture for the models on an FPGA. Our results show that Knowledge Distillation (KD) improves all tested U-Net variants. Among the evaluated platforms, the selected FPGA implementation achieves $398$ Frames Per Second (FPS) at $204.99$ Frames/J while maintaining a mean Intersection over Union (IoU) of $69.42\%$. In addition, our best model reaches $71.92\%$ mean IoU, which is $8.82$ percentage points (pps) higher than the previously reported result on the CrackVision12K dataset.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Inspection and maintenance of civil infrastructure, such as bridges, tunnels, and high-rise buildings, are essential to ensure long-term safety and reliability. Traditional inspection methods are largely manual, requiring human experts to access potentially dangerous locations to visually assess structural damage. With the emergence of UAVs, automated large-scale inspection of concrete and asphalt surfaces has become feasible, offering both safety and efficiency advantages. UAV-based inspection systems equipped with onboard computer vision can identify cracks and other forms of surface degradation without exposing human inspectors to hazardous environments.
The main challenge for onboard crack segmentation is the need to operate on resource-constrained edge devices with limited power, memory, and computational capacity, while still enabling real-time processing. Although Transformer-based models achieve state-of-the-art segmentation performance, their computational and memory demands hinder practical edge deployment. Consequently, lightweight architectures such as U-Net remain more suitable. Despite their simpler design, U-Net variants exhibit substantial variability in parameter count and computational cost, and still require optimization for efficient edge execution. Identifying models that balance segmentation accuracy and efficiency is therefore critical for real-world edge crack segmentation.
Efficient deployment of crack segmentation models on edge devices is governed by strict constraints on latency, power consumption, and energy efficiency, particularly under small batch sizes dictated by the number of input sensors (e.g., cameras). In this work, we investigate implementation across heterogeneous edge platforms, including CPUs, GPUs, and Neural Processing Units (NPUs), each offering distinct trade-offs. GPUs provide high throughput, especially for large batch sizes, but typically incur higher power consumption and are less efficient in small-batch regimes. CPUs offer greater flexibility and broader operator support, yet often fall short in terms of parallel efficiency. NPUs, while promising superior energy efficiency, are constrained by limited support for operators and model architectures, complicating deployment. These disparities make platform selection, numerical precision, and runtime configuration non-trivial design choices. Collectively, these limitations highlight the need for custom Field-Programmable Gate Array (FPGA)-based architectures to achieve superior energy efficiency and mitigate platform-specific constraints.
@@
  • We perform a comprehensive design space exploration across multiple levels, from model-level design to hardware platforms, providing a comparative analysis of segmentation results, latency, power consumption, and energy efficiency for crack segmentation on several edge platforms.
  • We design and implement U-Net variants spanning three orders of magnitude in parameter count to study the trade-off between segmentation performance and efficiency. Knowledge distillation from a Transformer model is applied to enhance the smaller U-Net variants. The resulting Pareto-optimal model is $63.63\times$ smaller, and 6.32pps better, compared to the currently best tested Hybrid-Segmentor model from Goo et al [11].
@@
  • The Pareto-optimal FPGA implementation outperforms alternative embedded platforms, achieving the highest energy efficiency (204.99 Frames/J) and throughput (398 FPS) while still improving the segmentation accuracy by 2.79pps compared to the baseline U-Net without KD.
@@ II-A Crack Segmentation Approaches
Early methods of crack segmentation are often based on graph-based approaches such as the use of shortest path algorithms, as in [12], or the use of minimum spanning trees (MST), as in [35]. The advantages of these algorithms lie in their efficient implementation and explainability. However, since the introduction of UNet [25], deep learning models recognize cracks much better. Due to its efficient encoder-decoder structure, skip connections, UNet has become particularly popular for crack segmentation, especially for real-time applications. As a result, there are now many deep learning approaches that build on this and propose new, improved models or training procedures [6, 20, 14, 19, 29, 18]. UNet++ was introduced in 2019 [34]. In 2023, an improved UNet incorporating an attention mechanism was proposed [1]. However, most of these variants require increased computational resources and are therefore unsuitable for edge or embedded platforms. One of the most advanced real-time capable models is UNext [2]. However, its reliance on attention mechanisms introduces architectural complexity that currently hinders edge deployments.
II-B Semantic Segmentation on Embedded Platforms
In recent years, embedded platforms have gained significant traction for crack segmentation. Some authors propose specifically designed segmentation models for embedded platforms. For example, [3] presents MCFF-L Net - an architecture with 1.18 million parameters - combined with knowledge distillation, which achieves good segmentation results and is able to run in real-time on a Jetson Xavier NX. Similarly, [33] considered crack segmentation on devices for structural health monitoring. A related approach by [32] uses a special KD approach for measurement tasks on embedded systems.
Other works focus on simplifying UNet for microcontroller-class hardware, like [8], which analyzed small CNN-based crack segmentation models on low-power microcontrollers. In contrast, [9] showed that UNet models can be compressed to meet the requirements of very resource-constrained devices such as the STM32 platform.
@@ II-C Semantic Segmentation on FPGA
Recently, FPGAs have become a popular choice for implementing semantic segmentation algorithms under real-time and limited power budget constraints. In [10], a compressed ENet is deployed entirely on-chip using HLS4ML, while [17] and [26] combine pruning, quantization, and Vitis-AI flow to run encoder-decoder networks in real time on FPGA. Other approaches redesign segmentation blocks specifically for hardware efficiency, such as [4]. Another approach to shrinking the model is about model compression and low-bit quantization. In [23], DeepLabV3+ is pruned using a hardware-aware genetic algorithm, while in [22] a UNet is quantized to 3-bit precision and implemented on a dedicated on-chip pipeline. Similar OpenCL- or HLS-based encoder-decoder accelerators further show the impact of quantization and optimized data flow on performance [31]. However, in the context of crack segmentation, FPGA-based solutions are less common. The work in [5] represents one of the few examples, implementing a particle-filter–based crack detector directly on an FPGA for real-time inspection. Although they use a totally different approach than we do, their approach already shows the relevance of FPGAs for building inspections.
@@ II-D FPGA-based U-Net Implementations
U-Net architectures pose additional challenges for streaming implementations due to their encoder–decoder structure with multiple skip connections that require buffering high-resolution feature maps. These skip connections substantially increase on-chip memory demand and can become the primary bottleneck in fully unrolled designs based on widely used FINN [28] and HLS4ML [7] hardware libraries. To address this limitation, we propose offloading skip connection feature maps to external memory. While this introduces additional memory accesses, the resulting access pattern is largely sequential and predictable, allowing efficient utilization of memory bandwidth. Consequently, the associated energy overhead is expected to be limited, while significantly relaxing on-chip memory constraints and enabling the deployment of larger U-Net variants.
@@ III Methodology
III-A Overview
Our objective is to enable efficient crack segmentation on resource-constrained edge devices, including embedded AI accelerators and FPGAs, without compromising accuracy. We systematically derive a family of U-Net variants by scaling the number of feature channels, creating models spanning three orders of magnitude in parameter count. To mitigate the performance degradation of smaller models, we leverage knowledge distillation from a high-capacity Transformer teacher, allowing compact U-Nets to retain segmentation accuracy close to the baseline. This approach, combined with hardware-aware optimization, facilitates the identification of Pareto-optimal models that balance efficiency and performance across diverse edge platforms.
III-B U-Net Channel Scaling
We use the original UNet model [25] as a basis and generate four smaller variants by reducing the number of feature maps. For this purpose, we introduce a parameter $c\in\{2,4,8,16,32\}$ which controls the number of all feature maps in the network while preserving the original encoder–decoder architecture. Setting $c=32$ reproduces the original UNet architecture. This parametrization allows us to analyze how the number of feature maps influences the segmentation results as well as the runtime and power consumption.
@@ III-B U-Net Channel Scaling
III-C Dataset and Preprocessing
In this work, we use the CrackVision12K dataset introduced by Goo et al. [11]. CrackVision12K is a curated benchmark for crack segmentation that combines images from several previously published crack datasets into a unified dataset with improved and standardized annotations. Compared with the original source datasets, the labels were manually refined to reduce annotation noise and inconsistencies. In addition, all images and masks were resized to a common spatial resolution of $256\times 256$ pixels.
In total, the dataset contains 12000 image-mask pairs. To ensure comparability with prior work on CrackVision12K, we use the train/validation/test split provided by Goo et al. [11]. The training set contains 9600 images, while the validation and test sets each contain 1200 images.
III-D Data Augmentation
@@ III-E Knowledge Distillation
III-F Training Setup and Hyperparameter Search
We first train all five U-Net variants by using supervised-learning to get the baseline models. In this setup, we use a combination of random search and an evolutionary algorithm as a hyperparameter search, adopted from Tschöpe et al. [27]. The optimized hyperparameters include the initial learning rate of the AdamW optimizer [21], the learning rate scheduler decay factor $\gamma$, the scheduler step size, the momentum values of the AdamW optimizer, the weight decay, and the class-weights that are used in the cross-entropy loss. However, for testing we use fixed class-weights, computed by the pixel distribution. We explain this in more detail in section III-G.
@@ III-F Training Setup and Hyperparameter Search
III-G Evaluation Metrics
We report both mean IoU and a weighted IoU. Let $\mathrm{IoU}_{bg}$ and $\mathrm{IoU}_{crack}$ denote the IoU values for the background and crack classes, respectively. The mean IoU is defined as
@@ III-G Evaluation Metrics
Thus, this metric weights the background pixels and crack pixels equally by a factor of $0.5$. Since crack pixels are much rarer but more important than background pixels in our application, we additionally report a weighted IoU:
$\mathrm{wIoU}=w_{bg}\cdot\mathrm{IoU}_{bg}+w_{crack}\cdot\mathrm{IoU}_{crack},$ (6)
@@
The weights are computed from the inverse class frequencies of the training set and normalized to sum to one. As a result, the weighted IoU emphasizes the crack class much more strongly than the background class. Therefore, a model can achieve a high mean IoU due to very good background segmentation while still obtaining a noticeably lower weighted IoU if its crack segmentation remains challenging.
@@ IV Implementation
This work investigates the performance of efficient U-Net implementations on a diverse range of embedded computing platforms. Besides FPGA, the comparison encompasses conventional CPU, GPU, and a dedicated AI accelerator, allowing for a comprehensive evaluation of deployment strategies under different computational and energy constraints.
@@ IV-A Edge Platforms
Specifically, we target: the Raspberry Pi 5 Model B (8 GB) featuring a quad-core Cortex‑A76 CPU, the Seeed Studio reComputer J3011 equipped with an NVIDIA Jetson Orin Nano (8 GB), the Coral Edge TPU M.2 module hosted on a Raspberry Pi 5. For CPU-based execution, the U-Net models were run using the ONNX Runtime with floating-point weights and activations. On the GPU platform, two inference configurations were explored: one fully relying on floating-point precision and the other employing int8 quantization for network parameters while retaining floating-point activations. The GPU experiments were conducted using the TensorRT execution engine to maximize efficiency. On the dedicated AI accelerator, the TensorFlow Lite models, quantized to int8 for both weights and activations, were compiled using the Edge TPU compiler. Inference latency was measured for batch size 1 and reported as a mean value of five execution runs on a subset of 1024 input images. The reported latency includes data transfer times between the host and the device, as well as the time required for conversions between fp32 and int8 data formats for inputs and outputs, respectively. Power consumption was monitored throughout the experiments using an external USB POWER‑Z KM003C KT002 multimeter. Both idle and runtime conditions were evaluated, with the latter involving continuous inference execution for a duration of 20 seconds. The mean power values were derived from three independent measurements. For the Coral Edge TPU setup, power readings correspond to the complete configuration, capturing the consumption of both the M.2 module and the Raspberry Pi 5 host. Active cooling solutions were employed for all Raspberry Pi devices, while the Coral Edge TPU operated with an aluminum heat sink to ensure thermal stability during execution.

### 06 May 2026 v2

@@ Abstract
Crack segmentation on edge devices can support continuous infrastructure monitoring and maintenance and thereby help to preserve public safety. Furthermore, autonomous infrastructure monitoring by using Unmanned Aerial Vehicles (UAVs) can reduce inspection risks, as human operators no longer need to enter hazardous areas. Edge processing reduces the cost of inspection by eliminating the need for high resolution image storage for offline processing and mitigates the security risks and bandwidth requirements of streaming to cloud servers. Edge inference is difficult due to the limited memory and computational capabilities of edge devices, which can affect both accuracy and latency. Furthermore, battery-powered devices are subject to strict power and energy constraints. Together, these limitations impose restrictions on the model size and computational complexity that can be deployed close to the sensor. Recent Transformer-based models reach high accuracy on semantic segmentation, but they are typically large and computationally intensive, which makes them hard to deploy on edge devices. We therefore use knowledge distillation to improve the segmentation results of smaller base models. We then use Post-Training Quantization (PTQ) to compress the models further. Additionally, we consider the deployment of these models across multiple edge platforms. To maximize energy efficiency, we design and implement a custom hardware architecture for the models on an FPGA. Our results show that Knowledge Distillation (KD) improves all tested U-Net variants. Among the evaluated platforms, the selected FPGA implementation achieves $398$ Frames Per Second (FPS) at $204.99$ Frames/J while maintaining a mean Intersection over Union (IoU) of $69.42\%$. In addition, our best model reaches $71.92\%$ mean IoU, which is $8.82$ percentage points (pps) higher than the previously reported result on the CrackVision12K dataset.
©2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Inspection and maintenance of civil infrastructure, such as bridges, tunnels, and high-rise buildings, are essential to ensure long-term safety and reliability. Traditional inspection is mostly manual: human experts have to access dangerous locations to look for structural damage. With the emergence of UAVs, automated large-scale inspection of concrete and asphalt surfaces has become feasible, offering both safety and efficiency advantages. UAV-based inspection systems equipped with onboard computer vision can identify cracks and other forms of surface degradation without exposing human inspectors to hazardous environments.
The main challenge for onboard crack segmentation is the need to operate on resource-constrained edge devices with limited power, memory, and computational capacity, while still enabling real-time processing. Transformer-based models reach high segmentation accuracy, but their compute and memory demands make them difficult to run on edge devices. Lightweight architectures like U-Net are therefore the more practical choice. However, U-Net variants differ a lot in parameter count and compute cost, and still need to be optimized for edge use. Identifying models that balance segmentation accuracy and efficiency is therefore critical for real-world edge crack segmentation.
Efficient deployment of crack segmentation models on edge devices is governed by strict constraints on latency, power consumption, and energy efficiency, particularly under small batch sizes dictated by the number of input sensors (e.g., cameras). In this work, we look at the implementation on different edge platforms, including CPUs, GPUs, and Neural Processing Units (NPUs), which all have different tradeoffs. GPUs provide high throughput, especially for large batch sizes, but typically incur higher power consumption and are less efficient in small-batch regimes. CPUs offer greater flexibility and broader operator support, yet often fall short in terms of parallel efficiency. NPUs, while promising superior energy efficiency, are constrained by limited support for operators and model architectures, complicating deployment. These disparities make platform selection, numerical precision, and runtime configuration non-trivial design choices. Collectively, these limitations highlight the need for custom Field-Programmable Gate Array (FPGA)-based architectures to achieve superior energy efficiency and mitigate platform-specific constraints.
@@
  • We explore the design space over model scaling, numerical precision, and hardware platform, and we compare segmentation accuracy, latency, power, and energy efficiency for crack segmentation on four edge platforms (CPU, GPU, TPU, and FPGA).
  • We design and implement U-Net variants spanning three orders of magnitude in parameter count to study the trade-off between segmentation performance and efficiency. We use KD from a Transformer model (teacher) to improve the smaller U-Net model (student). Under the train/val/test split released in [12] for CrackVision12K, our most accurate model reaches 8.82pps higher mean IoU than their Hybrid-Segmentor model, while using $30.64\times$ fewer parameters.
@@
  • Our Pareto-optimal FPGA implementation reaches the highest energy efficiency (204.99 Frames/J) and throughput (398 FPS) of all tested platforms, while also improving the mean IoU by 2.79pps over the baseline U-Net without KD.
@@ II-A Crack Segmentation Approaches
Early methods of crack segmentation are often based on graph-based approaches, such as shortest path algorithms [13] or minimum spanning trees [39]. These methods are efficient and explainable, but since the introduction of U-Net [28], deep learning models recognize cracks much better. Due to its efficient encoder-decoder structure with skip connections, U-Net has become particularly popular for crack segmentation, especially for real-time applications. As a result, there are now many deep learning approaches that build on this and propose new, improved models or training procedures [6, 21, 15, 20, 32, 19, 8]. U-Net++ was introduced in 2019 [38], and in 2023, an improved U-Net incorporating an attention mechanism was proposed [1]. However, most of these variants require increased computational resources and are therefore not suitable for edge platforms. One of the most advanced real-time capable models is UNext [2], but its reliance on attention mechanisms introduces architectural complexity that currently hinders edge deployments. More recent work explores Mamba- and state-space-model-based crack segmentation [22], which reduces the parameter count further. In our work, we instead start from a standard U-Net and recover the accuracy of compact variants via knowledge distillation, an approach that has also been shown to be effective for crack segmentation in [35].
II-B Semantic Segmentation on Edge Platforms
In recent years, edge platforms have gained significant traction for crack segmentation. Some authors propose specifically designed segmentation models for edge platforms. For example, [3] presents MCFF-L Net - an architecture with 1.18 million parameters - combined with knowledge distillation, which achieves good segmentation results and is able to run in real-time on a Jetson Xavier NX. Similarly, [37] considered crack segmentation on devices for structural health monitoring. A related approach by [36] uses a special KD approach for measurement tasks on edge systems.
Other works focus on simplifying U-Net for microcontroller-class hardware, like [9], which analyzed small CNN-based crack segmentation models on low-power microcontrollers. In contrast, [10] showed that U-Net models can be compressed to meet the requirements of very resource-constrained devices such as the STM32 platform.
@@ II-C Semantic Segmentation on FPGA
Recently, FPGAs have become a popular choice for implementing semantic segmentation algorithms under real-time and limited power budget constraints. In [11], a compressed ENet is deployed entirely on-chip using HLS4ML, while [18] and [29] combine pruning, quantization, and Vitis-AI flow to run encoder-decoder networks in real time on FPGA. Other approaches redesign segmentation blocks specifically for hardware efficiency, such as [4]. Another approach to shrinking the model is about model compression and low-bit quantization. In [26], DeepLabV3+ is pruned using a hardware-aware genetic algorithm, in [25] a U-Net is quantized down to 4-bit and implemented using a shift based accelerator, while in [24] a U-Net is quantized to 3-bit precision and implemented on a dedicated on-chip pipeline. Similar OpenCL- or HLS-based encoder-decoder accelerators further show the impact of quantization and optimized data flow on performance [34]. However, in the context of crack segmentation, FPGA-based solutions are less common. One of the few works in this direction is [5] , which implements a particle-filter-based crack detector directly on an FPGA for real-time inspection. Although they use a totally different approach than we do, their approach already shows the relevance of FPGAs for building inspections.
@@ II-D FPGA-based U-Net Implementations
U-Net architectures pose additional challenges for streaming implementations due to their encoder–decoder structure with multiple skip connections that require buffering high-resolution feature maps. These skip connections substantially increase on-chip memory demand and can become the primary bottleneck in fully unrolled designs based on widely used FINN [31] and HLS4ML [7] hardware libraries. To address this limitation, we propose offloading skip connection feature maps to external memory. This adds extra memory accesses, but the access pattern is mostly sequential and predictable, so we can use the memory bandwidth efficiently. We therefore expect only a small energy overhead, while we can free up enough on-chip memory to fit larger U-Net variants.
@@ III Methodology
III-A U-Net Channel Scaling
We use the original U-Net model [28] as a basis and generate four smaller variants by reducing the number of feature maps. For this purpose, we introduce a parameter $c\in\{2,4,8,16,32\}$ which controls the number of all feature maps in the network while preserving the original encoder–decoder architecture. Setting $c=32$ reproduces the original U-Net architecture. This parametrization allows us to analyze how the number of feature maps influences the segmentation results as well as the runtime and power consumption.
@@ III-A U-Net Channel Scaling
III-B Dataset and Preprocessing
In this work, we use the CrackVision12K dataset [12]. CrackVision12K is a curated benchmark for crack segmentation that combines images from several previously published crack datasets into a unified dataset with improved and standardized annotations. Compared with the original source datasets, the labels were manually refined to reduce annotation noise and inconsistencies. In addition, all images and masks were resized to a common spatial resolution of $256\times 256$ pixels.
In total, the dataset contains 12000 image-mask pairs. To ensure comparability with prior work on CrackVision12K, we use the train/validation/test split provided in [12]. The training set contains 9600 images, while the validation and test sets each contain 1200 images.
III-C Data Augmentation
@@ III-D Knowledge Distillation
III-E Training Setup and Hyperparameter Search
We first train all five U-Net variants with standard supervised learning to obtain the baselines. Then we optimize Hyperparameters with a combination of random search and an evolutionary algorithm, adopted from [30]. The search covers the AdamW [23] learning rate, momentum parameters, and weight decay, the scheduler decay $\gamma$ and step size, and the training-time class weights. For KD training, we additionally optimize the temperature $T$, the mixing factor $\alpha$, and the Dice weight $\lambda_{Dice}$. Training-time class weights are not used at evaluation. Instead, we use the fixed weights described in Section III-F for all reported numbers.
@@ III-E Training Setup and Hyperparameter Search
III-F Evaluation Metrics
Crack segmentation is in general a task with a highly imbalanced class distribution, as cracks are often very thin and small and occur much less frequently than the background. This is also the case with the CrackVision12K dataset that we use. If we only used the mean IoU as a metric, we would weight the background class and the much more interesting crack class equally, which would end in a deceptively high evaluation score. Therefore, in addition to mean IoU, we also use weighted IoU. Below, we present the definitions of both metrics in more detail.
Let $\mathrm{IoU}_{bg}$ and $\mathrm{IoU}_{crack}$ denote the IoU values for the background and crack classes, respectively. The mean IoU is then defined as:
@@ III-F Evaluation Metrics
The weighted IoU instead emphasizes the crack class more strongly than the background class:
$\mathrm{wIoU}=w_{bg}\cdot\mathrm{IoU}_{bg}+w_{crack}\cdot\mathrm{IoU}_{crack}$ (6)
@@
We compute the weights from the inverse class frequencies of the training set and normalize them so that they sum to one. The weights are fixed for all our experiments and are therefore not part of the hyperparameter search. A large gap between mean IoU and weighted IoU of a model means that most of its mean IoU comes from the background pixels and not from the crack pixels themselves.
@@ IV Implementation
In this paper we investigate the performance of U-Net implementations on four representative edge computing platforms: a conventional CPU, an edge GPU, a dedicated AI accelerator (TPU), and a custom FPGA design. The selected platforms cover a wide range of edge hardware under different computational and energy constraints.
@@ IV-A Edge Platforms
We consider three edge platforms: The Raspberry Pi 5 Model B (8 GB) with a quad-core Cortex-A76 CPU, the Seeed Studio reComputer J3011 with an NVIDIA Jetson Orin Nano (8 GB), and the Coral Edge TPU M.2 module hosted on a Raspberry Pi 5. On the CPU we run the U-Net via ONNX Runtime in fp32. On the GPU we use TensorRT in two configurations: full fp32, and int8-weights with fp32-activations. On the Edge TPU we deploy fully int8-quantized TensorFlow Lite models compiled with the Edge TPU compiler.
To measure inference latency, we use a batch size of 1 and calculate the average across five execution runs on a subset of 1,024 input images, including host–device transfer and fp32/int8 conversion overhead. Power is measured with an external USB POWER-Z KM003C KT002 multimeter, both at idle and during 20 s of continuous inference, averaged over three measurements. For the Coral Edge TPU, the measurement covers both the M.2 module and its Raspberry Pi 5 host. All Raspberry Pi devices use active cooling; the Edge TPU uses an aluminum heat sink.